OpenStudy (anonymous):

Using complete sentences, explain how you would use the quadratic formula to solve x^2 + 8x = –2. Why is the quadratic formula the best method to use? Using complete sentences, explain how you would use the quadratic formula to solve x^2 + 8x = –2. Why is the quadratic formula the best method to use? @Mathematics

OpenStudy (mathteacher1729):

Add two to both sides x^2 + 8x + 2 = 0 There is no easy way to factor this, because there are no two numbers which multiply to (2) yet add up to (8). The quadratic formula will give us the answers we seek. :)

OpenStudy (mathteacher1729):

CLARIFICATION because there are no two numbers which multiply to (2) yet add up to (8). There are no two INTEGERS which multiply to (2) yet add up to (8).

OpenStudy (anonymous):

The quadratic formula gives a general solution to any equation of the form: \[ax^2 + bx + c = 0 \] For any real value of a, b and c of our choice. So if solving the equation through factoring is too difficult the quadratic formula will always give the solution.

hero (hero):

x^2 + 8x + 2 = 0 x^2 + (4 + sqrt(14)x + (4 - sqrt(14)x + 2 = 0 x(x + 4 + sqrt(14) + (4-sqrt(14))(x + 4 + sqrt(14)) = 0 (x + 4 + sqrt(14))(x + 4 - sqrt(14)) = 0

hero (hero):

It's a mess, but it's right